A company estimates that it has 30% chance of being successful in bidding on a $45,000 contract. If it costs $8000 in consultant fees to prepare the bid, what is expected gain or loss for the company if it decides to bid on this contract?
Answer by Theo(13342)

the expected value is .30 * 45000 minus 1.00 * 8000.
this is because they gain 45,000 30% of the time but they lose 8,000 100% of the time because every contract has to be bid on whether they win the contract or lose it.

what does this mean?

obviously, it they win the contract,they make 45,000 and deduct 8,000 for a gain of 37,000.

if they lose the contract, they lose 8,000.

the expected value assumes many contracts of the same type are being bid on.

assume there were 1,000 contracts with each contract carrying the same probability.

they would win 30% of the contracts and they would lose 70% of the contracts.

since they bid on all of them, they would lose 1,000 * 8,000 = 8,000,000 dollars.
since they win 30% of them, they would gain 300 * 45,000 = 13,500,000 dollars.
their net gain would be 13,500,000 - 8,000,000 = 5,500,000.
divide that by 1,000 and their average net gain per contract bid on would be equal to 5,500.
this is the same as the expected value calculated above.